simplify: 6^6+6^6+6^6+6^6+6^6+6^6 (Leave your answer in the exponential form.)
Answer by stanbon(75887)

6^6+6^6+6^6+6^6+6^6+6^6
=6(6^6)=6^7
Cheers,
Stan H.